### v1

# The Effect of Language Diversity When Fine-Tuning Large Language Models for Translation

Abstract
Prior research diverges on language diversity in LLM fine-tuning: Some studies report benefits while others find no advantages. Through controlled fine-tuning experiments across 132 translation directions, we systematically resolve these disparities. We find that expanding language diversity during fine-tuning improves translation quality for both unsupervised and—surprisingly—supervised pairs, despite less diverse models being fine-tuned exclusively on these supervised pairs. However, benefits plateau or decrease beyond a certain diversity threshold. We show that increased language diversity creates more language-agnostic representations. These representational adaptations help explain the improved performance in models fine-tuned with greater diversity. [^1] We will release our code and models upon acceptance.
1 Introduction
General-purpose LLMs like Llama 3 (Grattafiori et al., 2024) show promise for machine translation but require targeted fine-tuning beyond their incidental bilingualism (Briakou et al., 2023) to match the performance of specialized translation systems. Through fine-tuning approaches ranging from two-stage methods (Li et al., 2024; Zeng et al., 2024; Stap et al., 2024) to more sophisticated optimization techniques (Xu et al., 2025; Zhu et al., 2024b), LLMs such as Tower (Alves et al., 2024) now outperform traditional NMT systems (Kocmi et al., 2024; Deutsch et al., 2025).
Current research presents conflicting evidence on multilingual fine-tuning strategies. Some studies show that scaling the number of tasks or languages during instruction tuning improves (cross-lingual) generalization (Wang et al., 2022; Muennighoff et al., 2023; Dang et al., 2024), while others report that just 1–3 fine-tuning languages effectively trigger cross-lingual transfer (Kew et al., 2024; Zhu et al., 2024a). Recent inference-only experiments by Richburg and Carpuat (2024) across 132 translation directions highlight this uncertainty, showing variance in translation quality with off-target generations for non-English sources and inconsistent performance across languages. While non-English over-tokenization and typological distance provide partial explanations, controlled fine-tuning experiments on the effects of language diversity during fine-tuning remain unexplored.
We address these conflicting findings through systematic experimentation with varying translation directions, measuring effects on both seen and unseen language pairs. Through controlled fine-tuning across 132 translation directions, we demonstrate that increasing language diversity consistently improves translation quality in all categories. Counterintuitively, models fine-tuned on the most diverse language sets outperform others even on fully supervised language pairs that less diverse models are specifically optimized to handle. However, experiments with even larger language sets (272 directions) reveal that benefits plateau or decrease beyond a certain diversity threshold. Analysis of model activations shows that fine-tuning on diverse language directions creates more target language-agnostic representations in middle layers, explaining the performance improvements in our most diverse models.
Figure 1: COMET-strict scores (left) and off-target rates (right) for base (no fine-tuning), fsec (English-centric), fs (seen directions), fs+ps+un (all directions), evaluated on fully supervised (de/en/ko/nl/ru/zh pairs), unsupervised (cs/is/ja/pl/sv/uk pairs), and partially supervised (combining supervised and unsupervised) language pairs. Numbers above bars show mean scores. Training on more diverse sets improves all categories, with fs+ps+un achieving best COMET-strict scores even for fully supervised pairs. fs substantially reduces off-target rates for unsupervised directions compared to base and fsec, despite these pairs being absent from its fine-tuning data.
2 Fine-tuning and evaluation design
Following Richburg and Carpuat (2024), we categorize our language pairs into three groups based on their presence in the fine-tuning data of the Tower model we build upon: fully supervised (pairs between de, en, ko, nl, ru and zh), zero-shot (pairs involving cs, is, ja, pl, sv and uk), and partially supervised (pairs combining supervised and zero-shot languages). This yields 132 translation directions across 12 typologically diverse languages with varying pre-training representation, enabling comprehensive assessment across different data conditions (see Appendix A).
Fine-tuning setups
We compare the following incremental fine-tuning approaches using the Tower family of models, which are built on Llama 2 and underwent continued pre-training with a mixture of monolingual and parallel data: base: TowerBase-7B model without task-specific fine-tuning, serving as our baseline.
fsec: base fine-tuned only on fully supervised English-centric translation directions (10 directions), representing minimal supervision.
fs: base fine-tuned on all fully supervised language directions (30 directions), extending beyond English-centric pairs to investigate transfer learning between diverse language combinations.
fs+ps+un: base fine-tuned on fully supervised, partially supervised, and unsupervised directions (132 directions), maximizing language diversity to investigate cross-lingual transfer effects.
This controlled experimental design allows us to systematically evaluate how increasing language diversity during fine-tuning affects both supervised and unsupervised translation directions, moving beyond aggregate scores to understand performance patterns across specific language groups.
Data
We fine-tune on NTREX-128 (Federmann et al., 2022), a high-quality dataset of 1,997 multi-parallel professionally translated sentences designed for machine translation evaluation. [^2] Preliminary experiments with additional FLORES-200 (dev) data showed no significant improvements, so we exclude it for experimental clarity. For evaluation, we use the FLORES-200 (Team et al., 2022) devtest set, which provides multi-parallel data for controlled cross-language comparison.
Metrics
Our primary metric is COMET-strict, a modified version of COMET (Rei et al., 2020) that assigns zero scores to off-target translations, following recommendations by Zouhar et al. (2024). [^3] We use version wmt22-comet-da. We also report off-target rates, measured using fastText (Joulin et al., 2017, 2016) language identification. [^4] We use the lid.176.bin model. Optimization and inference details are provided in Appendix B.
3 Results
Increased diversity leads to better performance
Figure 1 (left) demonstrates that expanding language diversity during fine-tuning yields consistent performance improvements across all language pair categories. The COMET-strict scores show a clear progression from base to fsec to fs to fs+ps+un models, with the most diverse model achieving the highest scores in every category. Surprisingly, the fs+ps+un model (fine-tuned on all 132 directions) outperforms specialized models even on fully supervised language pairs (0.880 vs. 0.876 for fsec), despite the latter being specifically optimized for these directions. The benefits become more pronounced for partially supervised (0.812 vs. 0.448 for base) and unsupervised (0.739 vs. 0.253 for base), although this improvement is expected as fs+ps+un is explicitly fine-tuned on these directions.
These results clarify conflicting evidence on language diversity (see §1) and align with Wang et al. (2022) and Dang et al. (2024), confirming that broad language diversity (132 directions vs. 10–30), rather than minimal exposure, substantially enhances cross-lingual transfer, even for pairs already well supported in more specialized models.
Increased diversity reduces off-target problem
Off-target translations, where models generate content in incorrect languages, represent a critical failure mode in LLM-based MT (Zhang et al., 2023; Guerreiro et al., 2023; Sennrich et al., 2024).
Figure 1 (right) shows that while all models maintain target language fidelity for fully supervised pairs, the base model produces incorrect target languages at alarming rates for partially supervised (44%) and unsupervised pairs (65%). Fine-tuning progressively mitigates this problem, with fs showing substantial improvements (13% and 31% respectively) despite not being explicitly trained on these language combinations. Significantly, the fs+ps+un model completely eliminates off-target translations across all categories.
Diversity benefits plateau
Expanding from fs+ps+un (132 directions) to 272 directions reveals nuances in the diversity-performance relationship. Unsupervised directions benefit from increased diversity, while fully supervised directions show slight performance decline, suggesting benefits plateau beyond a certain threshold (details in Appendix C.1). This contradicts prior work that found monotonic improvements with diversity Wang et al. (2022); Dang et al. (2024), but aligns with Muennighoff et al. (2023)’s observation of diminishing returns when scaling multilingual pretraining beyond certain language counts.
Regularization alone insufficient
Regularization benefits models by enhancing generalization and calibration, with strong effects when using distant languages (Meng and Monz, 2024). We investigate whether these benefits can be achieved through explicit regularization techniques (weight decay and LoRA) rather than language diversity, but find no comparable improvements. This aligns with Aharoni et al. (2019), who suggest that multilingualism provides benefits beyond explicit regularization methods. See Appendix C.2 for details.
Results not due to multi-parallel data
[…]
Findings persist at larger scale
Larger models (13B) exhibit the same trends: increased language diversity leads to reduced off-target rates and improved cross-lingual transfer. This confirms our findings are robust across model scales. Complete experimental details are provided in Appendix C.4.
Middle layers adapt most
Figure 2: SVCCA similarity scores between fine-tuned and base models across layers. Lower values indicate greater adaptation during fine-tuning. base-fsec (blue), base-fs (red), and base-fs+ps+un (green) are compared, with their mean shown in black. Shaded regions represent confidence intervals. Middle layers show most significant adaptation, with lowest mean similarity (0.91) at layer 12. fp+ps+un exhibits greater adaptation throughout the network.
We analyze activation patterns across models by comparing them with the base model using Singular Vector Canonical Correlation Analysis (SVCCA; Raghu et al., 2017). This analysis identifies where and to what extent adaptations occur during fine-tuning. We aggregate activations across all source-target language pairs and present the layer-specific results in Figure 2.
Our analysis reveals that middle layers consistently undergo the most substantial adaptation across all fine-tuned models, with the lowest mean similarity (0.91) occurring at layer 12. Furthermore, models fine-tuned on more languages exhibit greater divergence from the base model, with fs+ps+un showing most substantial adaptations.
Middle layers encode semantic information and show the strongest cross-lingual transfer capabilities (Liu and Niehues, 2025; Liu et al., 2025). Our findings support that larger degrees of cross-lingual transfer within middle layers explain the performance improvements observed in models fine-tuned on a larger linguistic diversity.
Diversity increases cross-lingual overlap
Figure 3: Kernel density estimation of layer 12 activations for base (top) and fs+ps+un (bottom). Colors represent translation directions. Intra-cluster distances show increased specialization for single-target clusters in fs+ps+un, while multi-target cluster C1 demonstrates increased overlap.
We analyze layer 12 (the most significantly modified layer) to understand which adaptations occur during fine-tuning. Following from Gao et al. (2024) and Wang et al. (2024), we apply t-SNE dimension reduction (van der Maaten and Hinton, 2008) to layer activations and visualize the bivariate kernel density (KDE) estimation. Next, we employ k-means clustering to identify language groups within these representations, using silhouette score maximization (Rousseeuw, 1987) for optimal cluster determination without requiring manual inspection. Finally, we calculate the intra-cluster distances. We compare the base and fs+ps+un models, visualizing unsupervised directions where we expect the most significant adaptations.
Figure 3 presents the resulting visualization. Notably, for the single-target language clusters C0 and C2, the fs+ps+un model exhibits greater intra-cluster distances (0.54±0.34 and 0.41±0.26) compared to the base model (0.45±0.32 and 0.35±0.20), suggesting increased specialization per source-target direction after fine-tuning on diverse data. Conversely, for the multi-target language cluster (C1), the fs+ps+un model shows reduced intra-cluster distances (0.47±0.29) relative to the base model (0.55±0.28), indicating greater representational overlap between these linguistically related languages. This increased overlap provides evidence for enhanced cross-lingual transfer, which contributes to the superior performance of models fine-tuned on greater linguistic diversity.
Table 1 presents intra-cluster distances for all models. Note that clusters contain the same languages for all setups. As diversity increases, single-target clusters (C0, C2) show greater specialization while multi-language cluster C1 exhibits enhanced representational overlap, suggesting improved cross-lingual transfer.
While previous work has explicitly aligned representations (Liu and Niehues, 2025; Kargaran et al., 2024; Stap et al., 2023), our findings show implicit alignment occurs through multilingual fine-tuning.
$\boldsymbol{\times}$ C0 $\boldsymbol{+}$ C1 $\boldsymbol{\star}$ C2
base $0.45\pm 0.32$ $0.55\pm 0.28$ $0.35\pm 0.20$
fsec $0.49\pm 0.33$ $0.53\pm 0.26$ $0.34\pm 0.20$
fs $0.52\pm 0.36$ $0.51\pm 0.28$ $0.39\pm 0.24$
fs+ps+un $0.54\pm 0.34$ $0.47\pm 0.29$ $0.41\pm 0.26$
Table 1: Intra-cluster distances. C0 (is target) and C2 (ja target) show increased distances in models fine-tuned on more diverse data, while C1 (cs, pl, sv, uk targets) shows decreased distances, indicating enhanced cross-lingual transfer.
4 Conclusion
Our systematic investigation across 132 translation directions resolves conflicting findings on language diversity in LLM fine-tuning. We show that fine-tuning on broader language sets consistently improves translation across all categories: fully supervised, partially supervised, and zero-shot pairs. Consequently, we recommend fine-tuning with diverse language directions even when optimizing for a limited subset of target translation pairs, as our most diverse model outperformed models specialized exclusively for those target pairs. However, we advise identifying an optimal diversity threshold, as too many languages diminishes performance for well-supported pairs while still benefiting less-represented languages. Our representational analysis attributes the diversity improvements to specific adaptations in middle layers, revealing increased language-agnostic representations, which explains the enhanced cross-lingual transfer.
Limitations
We evaluate on the FLORES-200 (Team et al., 2022) devtest set, a multi-parallel benchmark consisting of documents originally written in English and professionally translated into multiple languages. While this may introduce some translationese effects, the multi-parallel nature enables controlled comparison across language pairs.
Our findings are based on the Tower model family (Alves et al., 2024) (7B and 13B), built on LLaMA 2 (Touvron et al., 2023). Further research should verify whether these patterns generalize to other model architectures and even larger model sizes.
Appendix A Language details
The selection of languages shown in Table 2, following the language selection from Richburg and Carpuat (2024), enables evaluation across varied typological properties and scripts while providing a systematic comparison between supervised languages (seen during fine-tuning) and zero-shot languages that share linguistic features with the supervised set. The languages in the zero-shot set were chosen to represent both varying degrees of resource support in the pre-training data and to have relationships to languages in the supervised set through language family, typological properties, or orthography.
Appendix B Implementation details
B.1 Optimization
[…]
For all experiments, we performed full fine-tuning using the AdamW optimizer (Loshchilov and Hutter, 2019) with 5% warm-up percentage and trained for one epoch. Based on development set performance, we selected the optimal configuration: a cosine learning rate scheduler with batch size of 256 and learning rate of $2\times 10^{-5}$. We implemented our fine-tuning experiments using the Hugging Face transformers library (Wolf et al., 2020) with DeepSpeed (Rasley et al., 2020).
B.2 Inference
For both fine-tuning and zero-shot inference, we used the prompt template shown in Table 3. We mask out the prompt during fine-tuning. We employed greedy decoding (beam size 1) to balance computational efficiency with comprehensive evaluation across all translation directions.
Appendix C Additional results
C.1 Scaling diversity to 272 languages
To investigate whether further increasing language diversity yields additional benefits, we compared our most diverse model from the main experiments (fs+ps+un with 132 directions) to an even more diverse setup including 272 translation directions. While maintaining a similar distribution of language families as in our main experiments, we added five additional languages:
  • Germanic family: Danish (da, North Germanic) and Afrikaans (af, West Germanic)
  • Slavic family: Slovak (sk, West Slavic) and Bulgarian (bg, South Slavic)
  • East Asian languages: Vietnamese (vi, different writing system but shares vocabulary with Chinese)
This selection maintains balanced representation across language families while introducing controlled diversity within each family. All additional languages are represented in both NTREX and FLORES-200.
Importantly, we evaluate both models on the same set of languages and directions as used throughout the paper. The additional languages are only used during fine-tuning to increase diversity, allowing us to measure their impact on the original set of translation directions.
Figure 4: COMET-strict scores comparing models trained on 132 directions and 272 directions. Both are evaluated on the original test set with the same language pairs as used throughout the paper. Unsupervised directions show clearest benefits from increased diversity (+0.01), while fully supervised directions show a slight decrease (-0.003), suggesting potential diversity trade-offs.
Figures 4 shows the performance comparison between our 132-direction model (fs+ps+un) and the expanded 272-direction model.
For fully supervised pairs, we observe a slight performance decrease (-0.003 COMET-strict) when scaling to 272 directions. Partially supervised directions show almost identical performance (+0.001), while unsupervised directions demonstrate the clearest benefit (+0.01) from increased language diversity.
These results suggest that language diversity benefits may plateau or even slightly decline for already well-represented language pairs. The slight reduction in fully supervised performance may indicate a trade-off between focused optimization and broader generalization, where extremely high diversity can dilute the model’s effectiveness for specific well-represented languages. Nonetheless, the continued improvements for unsupervised directions (with respect to the original TowerBase model) demonstrate that higher diversity provides additional benefits for these previously unseen language combinations, even though both the 132 and 272 direction models include these pairs during fine-tuning.
C.2 Regularization alone insufficient
Figure 5: COMET-strict scores comparing fs models with weight decay values of 0.01 (standard) and 0.10. Increasing regularization strength shows minimal impact on fully supervised and partially supervised directions, while actually harming performance on unsupervised directions, suggesting that regularization alone cannot replicate the benefits of increased language diversity.
Figure 6: COMET-strict scores for 7B models fine-tuned on filtered NLLB dataset: base (no fine-tuning), fsec (English-centric), fs (seen directions), fs+ps+un (all directions), evaluated on fully supervised (de/en/ko/nl/ru/zh pairs), unsupervised (cs/is/ja/pl/sv/uk pairs), and partially supervised (combining supervised and unsupervised) language pairs. Numbers above bars show mean scores. Training on more diverse sets improves all categories, with fs+ps+un achieving best results even for fully supervised pairs. fs substantially reduces off-target rates for unsupervised directions compared to base and fsec, despite these pairs being absent from its fine-tuning data.
To investigate whether the performance benefits observed with increased language diversity could be achieved through explicit regularization techniques, we conduct additional experiments using stronger regularization on models with limited language diversity. If increased language diversity primarily functions as a form of regularization, we hypothesize that similar improvements could be obtained by directly increasing regularization strength in less diverse models.
All our previous experiments use the AdamW optimizer with weight decay set to 0.01 and gradient clipping at 1.0. This aligns with common practices in LLM fine-tuning, where dropout (Srivastava et al., 2014) is rarely employed (neither the Llama nor Tower papers mention dropout, though both use weight decay). Notably, AdamW applies weight decay directly to the weights rather than through gradients, decoupling it from the learning rate.
We tested this hypothesis by fine-tuning the fs setup with increased weight decay values of 0.05 and 0.10 (compared to our standard 0.01). We chose the fs setup to examine whether stronger regularization would induce better cross-lingual transfer to partially supervised and unsupervised directions, potentially mimicking the benefits observed in the more diverse fs+ps+un model.
Figure 5 shows the COMET-strict scores comparing fs models with weight decay values of 0.01 (standard) and 0.10.555The results for weight decay at 0.05 were very similar to 0.10 and are omitted for clarity. Increasing the regularization strength has minimal impact on translation performance across all language categories. For fully supervised directions, both models achieved identical mean scores (0.875). For partially supervised directions, the difference was negligible (0.699 vs. 0.697). For unsupervised directions, the model with stronger regularization actually performed slightly worse (0.483 vs. 0.490).
We further explored alternative regularization approaches by implementing LoRA (Hu et al., 2022) with rank 64, which constrains fine-tuning to a low-dimensional subspace. This parameter-efficient tuning method can be considered a form of regularization as it restricts model updates to a much smaller parameter space than full fine-tuning, potentially preventing overfitting. Results from LoRA experiments align with our weight decay findings: performance for fully and partially supervised directions remained comparable to full fine-tuning with standard regularization, while unsupervised directions showed slight degradation.
These experiments demonstrate that our initial weight decay value of 0.01 already provides an appropriate balance between overfitting prevention and model flexibility. More importantly, they confirm that the cross-lingual transfer benefits observed in more diverse models cannot be replicated merely by increasing explicit regularization in less diverse models. The language diversity benefits we observe go beyond simple explicit regularization effects, providing specialized cross-lingual knowledge transfer. Our findings align with Aharoni et al. (2019), who suggest that multilingualism provides benefits beyond what can be achieved through explicit regularization methods.
Figure 7: COMET-strict scores for 13B models: base (no fine-tuning), fsec (English-centric), fs (seen directions), fs+ps+un (all directions), evaluated on fully supervised (de/en/ko/nl/ru/zh pairs), unsupervised (cs/is/ja/pl/sv/uk pairs), and partially supervised (combining supervised and unsupervised) language pairs. Numbers above bars show mean scores. Training on more diverse sets improves all categories, with fs+ps+un achieving best results even for fully supervised pairs. fs substantially reduces off-target rates for unsupervised directions compared to base and fsec, despite these pairs being absent from its fine-tuning data.
C.3 Results not due to multi-parallel data
To verify our findings are not artifacts of using multi-parallel data, we constructed a non-multi-parallel dataset from the NLLB corpus (Team et al., 2022). We maintained the same 132 language directions as in our main experiments but eliminated the multi-parallel property Following Koehn (2024), we extract examples with LASER (Artetxe and Schwenk, 2019) scores above 1.05. We then removed sentences that appeared in multiple language pairs and sampled the remaining data to ensure exactly 2,000 examples per direction, creating a completely non-multi-parallel dataset of equivalent size to our NTREX experiments.
Figure 6 shows COMET-strict (left) and off-target (right) results from experiments conducted using the filtered NLLB dataset rather than NTREX, allowing us to verify that our findings are not artifacts of using multi-parallel data.
The results demonstrate that our core finding—increased language diversity during fine-tuning leads to better performance—holds when using non-multi-parallel data as well. The fs+ps+un model still achieves the highest COMET-strict scores across all language categories, including for fully supervised language pairs. This confirms that the benefits of diverse fine-tuning extend beyond the multi-parallel setting described in our main experiments.
When comparing performance between models fine-tuned on NLLB versus NTREX data, we observe identical ranking patterns across different fine-tuning setups, though the NTREX-trained models show slightly better overall performance. This marginal improvement is likely attributable to NTREX’s higher data quality, as it consists of professionally translated content specifically designed for machine translation evaluation.
C.4 Invariance to model scale
Figure 7 demonstrates that our findings about language diversity benefits persist when scaling to 13B parameters.
For translation quality (Figure 7, left), the most diverse setup (fs+ps+un) consistently achieves the best results across all language categories, including fully supervised pairs. While most 13B models show higher scores than their 7B counterparts (Figure 1, left), the fsec model unexpectedly performs worse than base in partially supervised and unsupervised settings (0.557 vs 0.558 and 0.154 vs 0.465), unlike in the 7B configuration where fsec outperformed base.
For off-target rates (Figure 7, right), the most diverse setup again eliminates off-target translations completely. No model produces off-target translations for fully supervised pairs. The fsec 13B model shows substantially worse performance for partially supervised (0.34) and unsupervised (0.80) pairs compared to its 7B version (Figure 1, right). Though base and fs 13B models show improved off-target rates compared to 7B, the problem remains significant (base: 39% for unsupervised, fs: 22%).
The decrease in performance for the fsec 13B model can likely be attributed to overfitting to the limited English-centric training data.
These results confirm that language diversity benefits during fine-tuning are robust across model scales, consistently improving both translation quality and target language fidelity.